# Topological classification of black holes surrounded by quintessence

### v1

@@ III Kerr black hole and it’s AdS case
The negative root $z_{3}$ does not satisfy the off-shell condition, so only the positive real roots are considered for the topological numbers. According to the residue theorem and the definition of the winding number equation (17), we have $w_{1}=1$, $w_{2}=-1$, resulting in a total topological number $W=0$ for the Kerr black hole in the presence of quintessence. This outcome is also inferred from the vector field $n$ depicted in Fig. 2(c). In other words, the winding and topological numbers for the Kerr black hole surrounded by quintessence match those of the standard Kerr case T5 , yet they deviate from the values of the Schwarzschild black hole under the same dark energy influence. Therefore, the spacetime rotation in the quintessence framework influences its topological characteristics.
For a special case where $\tau=\tau_{c}=\frac{12\sqrt{3}\pi a}{2-6\sqrt{3}ca}$, it is found that $\mathcal{A}(z)$ has two positive roots that coincide:
[…]
Denoting this coincident point as $z_{12}$, the characteristic function (27) can be explicitly written as
$\mathcal{R}(z)=-\frac{z^{2+3\omega_{q}}-a^{2}z^{3\omega_{q}}+3cz\omega_{q}}{(z-z_{12})^{2}(z-z_{3})}.$ (34)
Here, the obtained characteristic function $\mathcal{R}(z)$ has a second-order pole at $z_{12}$, which is a generation point. When the order $m$ of a pole $z_{i}$ satisfies $m\geq 2$, the residue of the analytic function $\mathcal{R}(z_{i})$ is given by
$\operatorname{Res}\mathcal{R}(z_{i})=\frac{1}{(m-1)!}\lim_{z\to z_{i}}\frac{d^{m-1}}{dz^{m-1}}\left[(z-z_{i})^{m}\mathcal{R}(z)\right].$ (35)
Thus, by calculating the residue of $\mathcal{R}(z)$ at the second-order pole $z_{12}$, the winding number at this special point is determined as
$w_{12}=-1.$ (36)
We observe that the winding number of this generation point differs from the results obtained using the $\phi$-mapping topological current theory and the spinodal curve calculation. As noted in Ref. Wei2 , the winding number of a generation point computed via topological current theory is 0, implying that the winding number of such points depends on the definition method. The same applies to annihilation points, as will be seen in the following discussion.
Next, we turn to the topological number of the Kerr-AdS black hole in quintessence. In Boyer-Lindquist coordinates, the spacetime of this black hole is given by kabh
@@
$T=\frac{\Delta r^{\prime}(r)}{4\pi(r^{2}+a^{2})}\bigg{|}{r=r_{+}}.$ (42)
@@
$\mathcal{F}=\frac{3(3+8\pi Pr_{+}^{2})(r_{+}^{2}+a^{2})-9cr_{+}^{1-3\omega q}}{6r+(3-8\pi Pa^{2})^{2}}-\frac{3\pi(r_{+}^{2}+a^{2})}{(3-8\pi Pa^{2})\tau}.$ (43)
@@
$\phi^{r_{+}}=$ $\frac{9(r_{+}^{2}-a^{2})+24\pi Pr_{+}^{2}(3r_{+}^{2}+a^{2})+27c\omega qr+^{1-3\omega q}}{2r+^{2}(3-8\pi Pa^{2})^{2}}$ (44)
@@
$\tau=\frac{4\pi r_{+}^{3+3\omega q}(3-8\pi Pa^{2})}{3r_{+}^{2+3\omega q}+24\pi Pr_{+}^{4+3\omega q}-3a^{2}r_{+}^{3\omega q}+8\pi Pa^{2}r_{+}^{2+3\omega_{q}}+9c\omega qr_{+}}.$ (46)
To determine the winding number, we define
$\mathcal{R}(z)=\frac{24\pi Pz^{4+3\omega_{q}}-3z^{3\omega_{q}}a^{2}+z^{2+3\omega_{q}}\left(3+8\pi Pa^{2}\right)+9cz\omega_{q}}{\tau\left(24\pi Pz^{4+3\omega_{q}}-3z^{3\omega_{q}}a^{2}+z^{2+3\omega_{q}}\left(3+8\pi Pa^{2}\right)+9cz\omega_{q}\right)+4\pi z^{3+3\omega_{q}}\left(-3+8\pi Pa^{2}\right)}.$ (47)
For $\omega_{q}=-2/3$, this simplifies to
$\mathcal{R}(z)=\frac{3z^{2}-6cz^{3}+24P\pi z^{4}+8P\pi a^{2}z^{2}-3a^{2}}{4\pi z^{3}\left(-3+8P\pi a^{2}\right)+\left(3z^{2}-6cz^{3}+24P\pi z^{4}+8P\pi a^{2}z^{2}-3a^{2}\right)\tau}.$ (48)
Taking the denominator of the equation, we define a polynomial function $\mathcal{A}(z)$ as
$\mathcal{A}(z)=$ $4\pi z^{3}\left(-3+8P\pi a^{2}\right)+\left(3z^{2}-6cz^{3}+24P\pi z^{4}\right.$ (49)
$\left.+8P\pi a^{2}z^{2}-3a^{2}\right)\tau.$
Figure 3(a) shows the variation of $r_{+}$ with $\tau$ where a generation point and an annihilation point (where $(d^{2}\tau/dr_{+}^{2})<0$) are marked with black dots. These points divide the black hole into three branches: a small black hole branch, an intermediate black hole branch, and a large black hole branch. To compute the topological numbers, we plot the roots of $\mathcal{A}(z)$ for three cases. For $\tau<\tau_{a}$ or $\tau>\tau_{b}$, Figure 3(b) shows the distribution of roots when $\mathcal{A}(z)=0$. In these cases, there is one positive real root and one negative real root, corresponding to a winding number and topological number of 1. Additionally, for $\tau_{a}<\tau<\tau_{b}$. In these cases, there is one positive real root and one negative real root, corresponding to a winding number and topological number of 1. Additionally, for
[…]
According to the residue theorem, the winding numbers for the positive real roots are $w_{1}=1$, $w_{2}=-1$, $w_{3}=1$, resulting in the same topological number $W=1$. This can also be seen in Fig. 3(c), which displays the unit vector field $n$ for $\tau_{a}<\tau<\tau_{b}$. Therefore, the Kerr–AdS black hole in a quintessence background exhibits a distinct topological number compared to Kerr black hole.
In Fig. 3(a), the generation and annihilation points are denoted by $\tau_{a}$ and $\tau_{b}$, respectively. The calculation of the topological number for $\tau=\tau_{a}$ using the residue theorem was discussed in the previous section and will not be repeated here. Below, we compute the topological number for $\tau=\tau_{b}$, where $z_{3}=z_{2}$
[…]
Denoting this coincident point as $z_{23}$, the characteristic function becomes
$\mathcal{R}(z)=\frac{3z^{2}-6cz^{3}+24P\pi z^{4}+8P\pi a^{2}z^{2}-3a^{2}}{(z-z_{1})(z-z_{23})^{2}(z-z_{4})}.$ (52)
Using the residue theorem, we calculate the residue of the characteristic function at its second-order pole $z_{23}$, yielding the topological number for this point:
$w_{23}=1.$ (53)
(a) The zero points of $\phi^{r_{+}}$ in the $\tau-r_{+}$ plane, a creation point and a annihilation located in $\tau/r_{0}=\tau_{a}/r_{0}=3.0101$ and $\tau/r_{0}=\tau_{b}/r_{0}=4.2354$.
(b) The roots of $\mathcal{A}(z)$.
(c) The unit vector field $n$ of $\tau/r_{0}=3.5$.
Figure 3: The topological properties of the Kerr-AdS black hole surrounded by quintessence, we set $a/r_{0}=0.1$, $cr_{0}^{2}=0.1$ and $Pr_{0}^{2}=0.1$.
IV Einstein-Gauss-Bonnet black hole
@@
It is evident that when $q=0$, the solution simplifies to the four-dimensional EGB black hole proposed by Glavan and Lin  Glavan . In the limit $\alpha\to 0$, the negative branch recovers the Schwarzschild black hole immersed in quintessence matter, as described in  Q4 . Taking both $q=0$ and $\alpha\to 0$, the metric reduces to the classical Schwarzschild solution. By adjusting the state parameter $\omega_{q}$, a variety of black hole configurations can be realized. For example, when $\omega_{q}=0$, , corresponding to the absence of quintessence, the metric again reduces to the standard 4D EGB black hole  Glavan . The function $f(r)$ and the location of the event horizon are highly sensitive to the parameters $\omega_{q}$, $q$, and the Gauss–Bonnet coupling $\alpha$. To ensure the physical validity of the quintessence field, $\omega_{q}$ must remain within the interval $-1<\omega_{q}<-1/3$, , which ensures a positive energy density of the dark energy component, given by  4D
@@
In equation (57), the "$\pm$" symbol represents the two branches of the solution. For large horizon radius $r$, equation (57) reduces to the four-dimensional Schwarzschild black hole solution surrounded by quintessence matter. Therefore, the negative branch of the solution is well-behaved. Consequently, we consider the negative branch solution in the following analysis, from which several thermodynamic quantities of the black hole can be derived as:
@@
Based on above thermodynamic quantities, the generalized off-shell free energy of the system can be derived
@@
$\phi^{r_{+}}=\frac{1}{2}+3q\omega_{q}r_{+}^{-1-3\omega_{q}}-\frac{2\pi(r_{+}^{2}+2\alpha)}{r_{+}\tau},$ (63)
@@
$\tau=\frac{4\pi r_{+}^{3\omega_{q}}(r_{+}^{2}+2\alpha)}{r_{+}^{1+3\omega_{q}}+6q\omega_{q}},$ (64)
In this case, the characteristic function is:
$\mathcal{R}(z)=$ $-\frac{z^{1+3\omega_{q}}+6q\omega_{q}}{4\pi z^{3\omega_{q}}(z^{2}+2\alpha)-\tau(z^{1+3\omega_{q}}+6q\omega_{q})}$ (65)
$=$ $-\frac{z^{1+3\omega_{q}}+6q\omega_{q}}{\mathcal{A}(z)}.$
Figure 4: The roots of $\mathcal{A}(z)=0$ for 4D EGB black hole. we set $\alpha=0.1$ and $q=0.1$
We analyze the roots of $\mathcal{A}(z)=0$. In order to examine the poles of the complex function $\mathcal{R}(z)$ and the corresponding winding numbers, we plot $\mathcal{A}(z)$ in Fig. 4. when $\tau_{c}<\tau$, the polynomial $\mathcal{A}(z)$ admits two positive real roots $z_{1}$ and $z_{2}$, satisfying $z_{1}<z_{2}$. Accordingly, the winding numbers are $w_{1}=1$ and $w_{2}=-1$, yielding a total topological number $W=0$. This result differs from the topological number of the Schwarzschild black hole under quintessence. Furthermore, Fig. 5 presents the zero-point diagram of $\phi^{r_{+}}$, from which it can be seen that the 4D EGB black hole features a generation point but lacks an annihilation point.
Figure 5: The zero points of $\phi^{r_{+}}$ in 4D and 5D EGB black holes are shown, where the black dots represent generation points. Let the parameters be $\alpha=0.1$ and $q=0.1$
For the case of 5D, the solution of this black hole is 5DQ
@@
The characteristic complex function in this case is
$\mathcal{R}(z)=$ $-\frac{z^{2+4\omega_{q}}+2q\omega_{q}}{2\pi z^{1+4\omega_{q}}(z^{2}+4\alpha)-\tau(z^{2+4\omega_{q}}+2q\omega_{q})}$ (73)
$=$ $-\frac{z^{2+4\omega_{q}}+2q\omega_{q}}{\mathcal{A}(z)}.$
@@
As shown in Fig. 5, we observe that the zeros of the vector field component $\phi^{r_{+}}$ for both 4D and 5D EGB black holes exhibit similar behavior. Therefore, the results for the 5D EGB black hole can be derived from those of the 4D EGB black hole. The findings indicate that for $\tau>\tau_{c}$, the winding numbers and the topological number are $w_{1}=1$, $w_{2}=-1$, and $W=0$, respectively.
Table 1: In the table below, we summarize the topological numbers of the black holes studied in this work, along with their generation points (GP) and annihilation points (AP). We also include results reported in some references for comparison.
Black hole spacetime Topological number GP AP
Schwarzschild BH Wei2 -1 0 0
Kerr BH T4 0 1 0
Kerr-AdS BH T4 1 1 or 0 1 or 0
Schwarzschild BH surrounded by quintessence -1 0 0
Kerr BH surrounded by quintessence 0 1 0
Kerr-AdS BH surrounded by quintessence 1 1 or 0 1 or 0
4D EGB BH surrounded by quintessence 0 1 0
5D EGB BH surrounded by quintessence 0 1 0
V Conclusions
This article explores the thermodynamic topological properties of black holes by utilizing the residue theorem. In this study, black hole solutions are regarded as defects in thermodynamic space, and a characteristic complex function is defined based on the generalized off-shell free energy. The singularities of this complex function are determined by the roots of the polynomial $\mathcal{A}(z)$. The winding number of each black hole branch is given by the residue of the characteristic function at its simple poles, which can be calculated using formula (17). The winding numbers reveal the local topological properties of the black hole system. By summing the winding numbers of all singular points, the global topological number is obtained, which aids in classifying different black hole spacetimes. Moreover, the winding numbers of generation points and annihilation points can be identified by the residues of the characteristic function at their second-order isolated singularities.
We studied the winding numbers and topological numbers of a family of black holes surrounded by quintessence, with the results summarized in Table I, demonstrating that the results obtained via the residue method agree with those calculated through the topological current theory. Within the chosen black hole models, we find that the Schwarzschild black hole in quintessence has a topological number of -1, with no generation or annihilation points. In contrast, the Kerr black hole and the 4D/5D Einstein-Gauss-Bonnet black holes under quintessence background have a topological number of 0, each featuring one generation point but no annihilation point. For the Kerr-AdS black hole in quintessence, its topological number matches that of the Kerr-AdS black hole without quintessence, with one generation point and one annihilation point identified. These results indicate that both the rotation parameter and the Gauss-Bonnet coupling parameter under the quintessence background alter the topological class of black holes.

### v2

@@ IV Schwarzschild-adS black hole surrounded by quintessence
Table 2: The orientation is indicated by the arrow of $\phi^{r_{+}}$ for the Sch-aBH-$\omega_{q}$ and Kerr-aBH-$\omega_{q}$ and the topological numbers..
BH spacetime $I_{1}$ $I_{2}$ $I_{3}$ $I_{4}$ $W$
Sch-aBH-$\omega_{q}$ $\to$ $\uparrow$ $\to$ $\downarrow$ 0
Kerr-aBH-$\omega_{q}$ $\to$ $\uparrow$ $\leftarrow$ $\downarrow$ 1
V Kerr-adS black hole surrounded by quintessence
Next, we turn to the topological class of the Kerr-anti de Sitter black hole surrounded by quintessence (Kerr-aBH-$\omega_{q}$). In Boyer-Lindquist coordinates, the spacetime of this black hole is given by kabh
@@ V Kerr-adS black hole surrounded by quintessence
$T=\frac{\Delta r^{\prime}(r)}{4\pi(r^{2}+a^{2})}\bigg|_{{r=r_{+}}}.$ (37)
@@
$\mathcal{F}=$ $\frac{3(3+8\pi Pr_{+}^{2})(r_{+}^{2}+a^{2})-9cr_{+}^{1-3\omega q}}{6r_{+}(3-8\pi Pa^{2})^{2}}$ (38)
$-\frac{3\pi(r_{+}^{2}+a^{2})}{(3-8\pi Pa^{2})\tau}.$
@@
$\phi^{r_{+}}=$ $\frac{9(r_{+}^{2}-a^{2})+24\pi Pr_{+}^{2}(3r_{+}^{2}+a^{2})+27c\omega_{q}r_{+}^{1-3\omega_{q}}}{2r_{+}^{2}(3-8\pi Pa^{2})^{2}}$ (39)
@@
$\beta=\frac{4\pi r_{+}^{3+3\omega_{q}}(3-8\pi Pa^{2})}{3r_{+}^{2+3\omega_{q}}+24\pi Pr_{+}^{4+3\omega_{q}}-3a^{2}r_{+}^{3\omega_{q}}+8\pi Pa^{2}r_{+}^{2+3\omega_{q}}+9c\omega_{q}r_{+}}.$ (41)
Figure 11 shows the variation of $r_{+}$ with $\beta$, a generation point and an annihilation point located in $\beta/r_{0}=\beta_{a}/r_{0}=3.0101$ and $\beta/r_{0}=\beta_{b}/r_{0}=4.2354$, respectively. These points divide the black hole into three branches: a small black hole branch, an intermediate black hole branch, and a large black hole branch.
Figure 11: The zero points of $\phi^{r_{+}}$ in the $r_{+}-\beta$ plane for the Kerr-aBH-$\omega_{q}$, we set $a/r_{0}=0.1$, $cr_{0}^{2}=0.1$ and $Pr_{0}^{2}=0.1$.
The characteristics of the vector field $n$ shown in Fig. 12, where three zero points, indicated by black dots, are identified. These zeros partition the black hole configurations into small, intermediate, and large types. Fig. 13 presents the evolution of the vector field components ($\phi^{r_{+}},\phi^{\Theta}$) traced along the closed contours $C_{i}$ depicted in Fig. 12. The zero point of $\Phi_{i}$ is located at the origin. Contours with three zeros (from left to right) possess winding numbers +1, -1 and +1, respectively. The asymptotic behavior of $\phi$ along the boundary segments $I_{i}$ is summarized in Table II: at $I_{1}$ the vector field points rightward, while at $I_{3}$ it points leftward, revealing its topological number of +1. In the regime of low temperature ($\beta\to\infty$), the configuration gives rise to stable small black holes; conversely, when $\beta\to 0$, stable large black holes dominate the system. Thus, the Kerr-aBH-$\omega_{q}$ is classified within the topological class $W^{1+}$.
Figure 12: The unit vector field $n$ on the $r_{+}-\Theta$ plane for the Kerr-aBH-$\omega_{q}$, we set $\beta/r_{0}=3.5$.
Figure 13: Contours $\Phi_{6}$, $\Phi_{7}$ and $\Phi_{8}$ depict the variation of the vector field $\phi$ components along the pats $C_{6}$, $C_{7}$ and $C_{8}$ show in Fig. 12 for the Kerr-aBH-$\omega_{q}$.
VI Einstein-Gauss-Bonnet black hole surrounded by quintessence
@@
It is evident that when $q=0$, the solution simplifies to the four-dimensional EGB black hole proposed by Glavan and Lin Glavan . In the limit $\alpha\to 0$, the negative branch recovers the Schwarzschild black hole immersed in quintessence matter, as described in Q4 . Taking both $q=0$ and $\alpha\to 0$, the metric reduces to the classical Schwarzschild solution. By adjusting the state parameter $\omega_{q}$, a variety of black hole configurations can be realized. For example, when $\omega_{q}=0$, corresponding to the absence of quintessence, the metric again reduces to the standard 4D EGB black hole Glavan . The function $f(r)$ and the location of the event horizon are highly sensitive to the parameters $\omega_{q}$, $q$, and the Gauss–Bonnet coupling $\alpha$. To ensure the physical validity of the quintessence field, $\omega_{q}$ must remain within the interval $1<\omega_{q}<-1/3$, , which ensures a positive energy density of the dark energy component, given by 4D
@@
In equation (45), the $\pm$” symbol represents the two branches of the solution. For large horizon radius $r$, equation (45) reduces to the four-dimensional Schwarzschild black hole solution surrounded by quintessence matter. Therefore, the negative branch of the solution is well-behaved. Consequently, we consider the negative branch solution in the following analysis, from which several thermodynamic quantities of the black hole can be derived as:
@@
here $A=4\pi r_{+}^{2}$ represents area of the black hole horizon. Based on above thermodynamic quantities, the generalized off-shell free energy of the system can be derived
@@
$\phi^{r_{+}}=\frac{1}{2}+3q\omega_{q}r_{+}^{-1-3\omega_{q}}-\frac{2\pi(r_{+}^{2}+2\alpha)}{r_{+}\tau}.$ (51)
@@
$\beta=\frac{4\pi r_{+}^{3\omega_{q}}(r_{+}^{2}+2\alpha)}{r_{+}^{1+3\omega_{q}}+6q\omega_{q}}.$ (52)
Figure 14 presents the zero-point diagram of $\phi^{r_{+}}$, from which it can be seen that the 4D EGB black hole surrounded by quintessence (4D-EGB-BH-$\omega_{q}$) features a generation point $\beta_{c}$ but lacks an annihilation point. Figures 15 and 16 indicate that for $\beta>\beta_{c}$, the winding numbers of the zero points enclosed by the contours $C_{9}$ and $C_{10}$ are 1 and -1, respectively, yielding a total topological number $W=0$. This outcome differs from that of the Schwarzschild black hole surrounded by quintessence matter. Nevertheless, it can be inferred that the 4D EGB black hole and the Kerr black hole in the same matter field belong to the identical universal thermodynamic class, $W^{0+}$. Furthermore, the asymptotic behavior of the vector field near the boundaries $I_{i}$ is summarized in Table III.
Figure 14: The zero points of $\phi^{r_{+}}$ in the $r_{+}-\beta$ plane for 4D/5D-EGB-BH-$\omega_{q}$, we set $\alpha/r_{0}^{2}=0.1$ and $qr_{0}=0.1$.
Figure 15: The unit vector field $n$ on the ${r_{+}}-\Theta$ plane for the 4D-EGB-BH-$\omega_{q}$, we set $\beta/r_{0}=20$.
Figure 16: Contours $\Phi_{9}$ and $\Phi_{10}$ depict the variation of the vector field $\phi$ components along the pats $C_{9}$ and $C_{10}$ show in Fig. 15 for the 4D-EGB-BH-$\omega_{q}$ case.
For the case of 5D EGB black hole surrounded by quintessence (5D-EGB-BH-$\omega_{q}$), the solution of this black hole is 5DQ
@@
As shown in Fig. 14, we observe that the zeros of the vector field component $\phi^{r_{+}}$ for both 4D and 5D EGB black holes exhibit similar behavior. The results for the 5D EGB black hole can be derived from those of the 4D EGB black hole. Therefore, the 5D-EGB-BH-$\omega_{q}$ belongs to the specific topological class, $W^{0-}$.
Table 3: The orientation is indicated by the arrow of $\phi^{r_{+}}$ for the 4D/5D-EGB-BH-$\omega_{q}$ and the topological numbers..
BH spacetime $I_{1}$ $I_{2}$ $I_{3}$ $I_{4}$ $W$
4D-EGB-BH-$\omega_{q}$ $\leftarrow$ $\uparrow$ $\leftarrow$ $\downarrow$ 0
5D-EGB-BH-$\omega_{q}$ $\leftarrow$ $\uparrow$ $\leftarrow$ $\downarrow$ 0
Table 4: In the table below, we summarize the universal thermodynamic topological classifications of the black holes surrounded by quintessence, together with a summary of their corresponding thermodynamic behaviors and their generation points (GP) and annihilation points (AP).
Black hole spacetime W Classes Innermost Outermos Low T High T DP
Sch-BH T14 $W^{1-}$ unstable unstable unstable large unstable small in pairs
Kerr-BH T17 $W^{0+}$ stable unstable unstable large + stable small no one more GP
Sch-aBH T14 $W^{0-}$ unstable stable no unstable small + stable large one more AP
Sch-BH-$\omega_{q}$ $W^{1-}$ unstable unstable unstable large unstable small in pairs
Kerr-BH-$\omega_{q}$ $W^{0+}$ stable unstable stable small + unstable large no one more GP
Sch-aBH-$\omega_{q}$ $W^{0-}$ unstable stable no unstable small + stable large one more AP
Kerr-aBH-$\omega_{q}$ $W^{1+}$ stable stable stable small stable large in pair
4D-EGB-BH-$\omega_{q}$ $W^{0+}$ stable unstable unstable large + stable small no one more GP
5D-EGB-BH-$\omega_{q}$ $W^{0+}$ stable unstable unstable large + stable small no one more GP
VII Conclusions
In this work, we have systematically investigated the universal thermodynamic topological classifications of a family of black holes surrounded by quintessence, including Schwarzschild, Kerr, AdS, and Einstein–Gauss–Bonnet (EGB) spacetimes. The results are summarized in Table IV. Our analysis reveals that the Schwarzschild black hole in a quintessence background belongs to the $W^{1-}$ class, showing thermodynamic instability in both low- and high-temperature regimes. The introduction of rotation fundamentally modifies the topological characteristics: the Kerr black hole transitions to the $W^{0+}$ class, allowing coexistence of a stable small and an unstable large black hole at low temperatures, while no stable configurations persist at high temperatures. The presence of a quintessence field does not alter this universal classification.
For the AdS counterparts, the thermodynamic topology becomes markedly distinct. The Schwarzschild–AdS black hole corresponds to the $W^{0-}$ class, featuring the emergence of an unstable small and a stable large black hole only at high temperatures, with no black hole solution at low temperatures. In contrast, the Kerr–AdS configuration is classified as $W^{1+}$, where both small and large black holes can maintain stability across temperature ranges. This distinction highlights the significant influence of the AdS curvature and rotational degrees of freedom on the thermodynamic topology.
Furthermore, the inclusion of higher-order curvature corrections through the Gauss–Bonnet term in both 4D and 5D spacetimes preserves the $W^{0+}$ classification. These cases exhibit similar thermodynamic structures to the Kerr black hole in quintessence, characterized by the coexistence of stable and unstable branches at low temperatures and the disappearance of black hole states at high temperatures.
Overall, the comparative analysis demonstrates that the topological classification of black holes is sensitive to intrinsic parameters such as rotation and higher-curvature couplings, while the quintessence field mainly shifts the temperature-dependent stability ranges without altering the topological class itself. This work thus provides a new perspective for understanding the global stability structures of black holes in dark energy backgrounds.